OpenStudy (anonymous):

two planes leave an airport at the same time. One flies north at 505 mph, the other flies south at 565 mph. How far apart will the two planes be in 3 hours

OpenStudy (anonymous):

The first plane is travelling at 505 mph, so in three hours, the plane wil travel 1515 miles. 505 * 3 = 1515 miles The second plane is travelling at 565 mph, so in three hours, the plane will travel 1695 miles. 565 * 3 = 1695 miles The planes will be 3210 miles apart, since: 1515 + 1695 = 3210 miles